A bullet of weight 50 m is fired into a body of weight 5 kg, which is suspended by a string 1 m long. Due to this impact, the 5 kg. body swings through an angle of 30 deg from the vertical. Calculate the velocity of the bullet.

Grade:12th pass

1 Answers

Baijnath
5 years ago
Since bullet is moving with some velocity...so it has KE=1/2mv^2
And it displace bigger block . .thr throan angle of 30° …
Draw fugfig...
U will fing that.....PE of block is... changed...i.e Mgh...
(With hight  L(1-cos©) )
So equate...KE And PE...
i.e 1/2mv^2=MgL(1-cos©)......(.1)
U will get the answer
Hope you understand 😊😊😊😊😊😊😊